Arlene wishes to invest $5000. if she invests part at 7% simple interest, part at 6%, and recieves a total of $332 after one year, how much does she invest at each rate?
Answer by stanbon(48535)

Arlene wishes to invest $5000. if she invests part at 7% simple interest, part at 6%, and recieves a total of $332 after one year, how much does she invest at each rate?
--------------
Let amount invested at 7% be "x" dollars.
Interest on that money is 0.07x dollars.
Amount invested at 6% is "5000-x" dollars.
Interest on that money is 0.06(5000-x)= 300-0.06x dollars
EQUATION:
interest + interest = 332 dollars
0.07x + 300-0.06x = 332
0.01x=32
x=$3200 (amount invested at 7%)
5000-x= $2800 (amount invested at 6%)
Cheers,
Stan H.